OpenStudy (anonymous):

the lengths of the sides of a triangle are 3, 3,and . Can the triangle be a right triangle?

OpenStudy (anonymous):

\[ and,3\sqrt{2}\]

OpenStudy (cwrw238):



OpenStudy (cwrw238):

hypotenuse = sqrt(3^2 + 3^2) = sqrt18 this should help

OpenStudy (anonymous):

yes, thank you

OpenStudy (cwrw238):

yw